OpenStudy (anonymous):

determine the integral: 1/square root of 3x - 7

OpenStudy (amistre64):

need a 2 for the bottom and a 3 for the top; multiply by 6/6

OpenStudy (amistre64):

\[\int\frac{2}{2}*\frac{3}{3}\frac{1}{\sqrt{3x-7}}\] \[\frac{2}{3}\int\frac{3}{2\sqrt{3x-7}}\]

OpenStudy (amistre64):

do you see why?

OpenStudy (anonymous):

\[\int\limits \frac{dx}{\sqrt{3x-7}}\] Let: \[u=\sqrt{3x-7} \rightarrow u^2=3x-7\] Then: \[2 u du=3dx \rightarrow \frac{2}{3}udu=dx\] Pluggin into the integral it becomes: \[\frac{2}{3} \int\limits \frac{u du}{u}=\frac{2}{3} \int\limits du=\frac{2}{3}u+C=\frac{2}{3}\sqrt{3x-7}+C\]

OpenStudy (amistre64):

lol .... you used a "u" :)

OpenStudy (anonymous):

xP

OpenStudy (anonymous):

You can also do u=3x-7. I just wasn't paying attention and did it the harder way.

OpenStudy (amistre64):

i just multiplied by 1 ;)

OpenStudy (anonymous):

I mean, yeah, but either way some people can see it off that :P

OpenStudy (anonymous):

I'm not doubting your method amistre :P Even though I'M STILL NOT A MOD :'(

OpenStudy (amistre64):

\[\color{purple}{\small\text{maybe tomorrow ?}}\]

OpenStudy (anonymous):

Please :D

OpenStudy (amistre64):

the position comes with a lantern and a ring; and a box of hohos

OpenStudy (amistre64):

lantern ring and hohos sold seperately

OpenStudy (anonymous):

I would be happy to accept the responsibility :)

OpenStudy (amistre64):

when everyones a mod; then noones a mod ;)

OpenStudy (anonymous):

But but butttttt I do alot on here :D I've only been here like 2 weeks and I have like 240 medals or something O:

OpenStudy (amistre64):

ive been here since mid march i think ....

OpenStudy (anonymous):

Week, feels like a year...